# A Survey of Language-Based Communication in Robotics

### v1

@@ 3.4.3. Dynamic Adaptation with Language
Some systems use language as a feedback mechanism with which to adjust an existing model. For example, a functional policy for a robot to walk can take in a textual description of the environment and can use this to modify the behaviour accordingly (Shek et al., 2023). This provides a promising means for robots to adjust to unseen conditions through the generality of language models. By invoking the LLM to check each step of a plan, iterative adaptation can be achieved within the model’s architecture. This can include checking each step of the plan one-by-one (Raman et al., 2022), or calling on external observations from other sources to update, improve, or complete each step of the plan (Xu et al., 2023b). Textual feedback can also be be used to catch mistakes when a robot’s plan is converted to a Planning Domain Definition Language (PDDL) and the plan is symbolically checked; error messages can be passed back to the model for the plan to be adjusted accordingly (Shirai et al., 2023). This same approach can also be used for parsing error messages from robot execution, which in turn are used to edit generated code in the same way a human might respond to error messages while programming (Skreta et al., 2023), or simply assessments of why a plan did not work which are then converted into summaries and revised plans (Shinn et al., 2024). Generally speaking, these methods use the environment (whether that be a physical environment, or a human) as a means to evaluate their solution and then iteratively improve the output though this process (Yao et al., 2023; Wu et al., 2023c). Another approach is to generate a language plan step by step, but to execute the plan as the steps are generated. After a step is completed, an environment update is fed back into the model so it receives real-time feedback and can adjust accordingly (Jin et al., 2023). Adaptation can also be used to adapt a shared plan iteratively for different goals or to resolve different issues, for example a “brain LLM” can write a high level plan which is adjusted by a “body LLM” to make it performable by the robot. This allows a single plan to be written, improved, and ultimately realised by successive language calls (Bhat et al., 2024).
4. Applications
@@ 4.1. Autonomous Vehicles
Although human drivers do not generally have the ability to communicate verbally with each other, some amount of implicit communication is used (pointing, flashing lights, etc). It is possible that the integration of LLMs will provide agents an interface with which to perform the complex coordination required on roam environments with a wide variety of vehicles each pursuing different end goals. The changeable nature of road systems such as contraflows, breakdowns, and weather are very difficult to train for with existing approaches, so LLMs may provide a useful means to interpret and handle unexpected scenarios (Cui et al., 2024b). A few works are actively pursuing LLMs as a control method for vehicles, taking in textual descriptions and sometimes images, and outputting a text-based command for how the wheel and pedals should be adjusted (Cui et al., 2023a, 2024a; Xu et al., 2023c). A key advantage of using LLMs in this way is that they can provide reasoning and justification, either as an observational tool to explain the actions of an existing model (Park et al., 2024), or to pair a reason with the predicted action of the human, which in turn could be used to control the vehicle (Wen et al., 2023). The poor ability to reason about the environment, along with obvious safety concerns likely render a fully integrated text-based self-driving car unlikely. Some have addressed this with hybrid approaches, combing the LLM with a more reliable rule-based planner (Sharan et al., 2023). However, the most practical avenue for LLMs in this domain is language as a tool for explanation (robot-to-human), control at a very high level (human-to-robot), and inter-agent negotiation e.g. at a stop junction (robot-to-robot). Such negotiation is typical of LLMs, whether in a competitive setting (Wu et al., 2024c), or in a more cooperative system (Hunt et al., 2024). The explainability problem is especially well-matched to autonomous vehicles, and is being actively explored (Tekkesinoglu et al., 2024); language is an especially promising avenue in this space, and naturally extends the way humans might explain their intentions and actions while driving (Yuan et al., 2024b).
4.2. Language-based Explainability of Robot Behaviour
@@
Explainability for language models is a popular area of research with a wide variety of sub-areas; the literature is still young and LLMs are still fundamentally a black-box technology that is difficult to understand (Zhao et al., 2024). Despite this, language is a natural domain for explanation, especially for non-technical users. LLMs are themselves used to explain the behaviour of other models (González-Santamarta et al., 2023), offering the additional advantage of open dialogue where a human can ask specific targetted questions of the model (Lakkaraju et al., 2022). Robotics has an especially high demand for explainability, as robots tend to be mistrusted in general (Lewis et al., 2018), and can often cause damage or loss of life if they malfunction (Setchi et al., 2020). The union of language and robotics creates an area with an absolute need for explainability and transparency in the absence of trustworthy execution. Robots that use language would be able to explain their failures (Das et al., 2021), or could continuously update operators and surrounding humans of their reasoning, intentions, and beliefs. This type of explainability may prove crucial to providing safe and trustworthy AI systems in the future.
4.3. Ad-hoc coordination
@@
As robots and other autonomous systems become more prevalent it will be necessary for teams of such devices to dynamically form and cooperate (Stone et al., 2010). For example a number of home robots may need to organise a group strategy to clean a home, a team of agricultural robots may need to combine efforts to expedite a task before the weather turns, or set of helper robots might even need to collaborate to save a human in danger. Each of these requires a coordination and control structure that is agnostic to the models and morphology of the platforms involved (Flavin and Sen, 2024). Language is already used by humans in this way, and is an important aspect of effective ad-hoc teams (White et al., 2018; Mirsky et al., 2020). Collaborative emergent behaviour can be achieved through the use of language (Li et al., 2023a). LLM based structures may also provide a means for differences in robot ability to be considered and optimised for during allocation and planning (Han et al., 2024b). In these types of systems a variety of possible topologies may be required featuring peer-to-peer or hierarchical structures, each requiring an organisational system which could benefit from language as a communication medium in the same way humans operate (Guo et al., 2024). In order to establish a shared understanding, robots may need to agree upon names and labels for objects in the environment based on sensor observations, after which language can be used to represent the complex details of this specific entity when coordinating towards a goal (Ekila et al., 2024). Finally, this type of coordination naturally extends to human-robot teams (Liu et al., 2023c; Gong et al., 2023).
4.4. Cobots
@@
The popular emerging model for robots in the workplace is the Collaborate Robot (“cobot”) model. This means that humans are not replaced with machines, but instead humans and robots work alongside each other. Many governments are pursuing this model to improve their local industries, especially for small-to-medium sized businesses without the need or budget for large and specialised equipment (Lefranc et al., 2022). A key advantage of this is that systems and processes could feasibility be structured to utilise the best of each party’s skills and abilities. Much research and many robot platforms have been applied to this problem (Taesi et al., 2023). Use cases of real deployment have been performed with reasonable success (El Makrini et al., 2018). Many different approaches have been proposed ranging from working simultaneously on the same widget to forming a human-robot production line with sequential work (El Zaatari et al., 2019). Regardless, this means of interacting with robots will likely soon be commonplace in not just the factory setting, but service and medical too (Djuric et al., 2016). Substantial trust issues exist due to the dangerous nature of industrial robots (Kopp et al., 2022) which may be abated somewhat through the careful use of language. For humans to effectively work alongside robots there must be a shared understanding and means of communication, at least in one direction as a means to task the robot (human-to-robot) (Bergman et al., 2019), and language is likely the most powerful and natural approach (Ionescu and Schlund, 2021). The inclusion of language in these robots’ policies in any of the ways described in this work could substantially resolve trust issues and provide an effective means for humans and robots to cooperate in this setting.
5. Challenges
@@
There are still many challenges associated with Language Models, both generally and in robotics. In this section we discuss some of the primary areas of concern.
5.1. Core Issues with LLMs
@@ 5.1.1. Hallucination
The term “hallucination” refers to foundational models inventing or imagining new information that has nothing to do with the inputs or internal knowledge. This problem is well documented and has persisted despite strong concerns that this may render the models almost unusable (Rawte et al., 2023). It has been suggested that the fundamental architecture and approach of these models, hallucination is impossible to fully eliminate (Xu et al., 2024). The attention has turned to mitigating this problem, primarily through detection and resolution, although approaches are varied (Tonmoy et al., 2024). The exact reasons behind hallucination are still not well understood, although some have attempted to analyse and find traces of the problem by investigating hidden states in the model (Duan et al., 2024). Some have proposed calling in additional knowledge to resolve the issue (Liang et al., 2024), or by considering the likelihood of the generated text (for example in vision-language models, a bird near a tree is a sensible output so is more trustworthy) (Zhou et al., 2023). Another approach which is relevant to the robot-to-robot category is to collaborate with other agents, with agents abstaining from a group decision if required (Feng et al., 2024). The most popular area of development is to detect hallucination so the outputs of the model can at least be ignored (Chen et al., 2023a; Leiser et al., 2024; Zhu et al., 2024b). This is essential as, like the closely related adversarial examples in image recognition (Wiyatno et al., 2019), hallucination events generally give highly erroneous results without a correspondingly low confidence; the model doesn’t know it is ignorant. This has been partially addressed by pairing the outputs of the model with an associated uncertainty value (Kirchhof et al., 2024; Han et al., 2024a).
@@
Hallucination is an issue in all LLM applications, although in robotics it may have more impactful negative consequences. An imagined danger which does not exist or an imagined empty space where there is in fact a human could result in a robot causing considerable harm or loss of life. Evidently for an embodied robot to be deployed in the real world, issues surrounding hallucination must be addressed. At the absolute least, such events should reliably be detected so the robot does not misclassified a human, or attempt to act upon an environment it does not truly understand.
5.1.2. Security
@@
Any control system for a robot can be subjected to cyber attacks. Hostile actors may seek to gain control of or to manipulate the robot in order to disrupt industrial processes or ransom a company by interrupting its operations. Connecting an LLM to one of these systems makes it a target. This is especially true with the increasing ubiquity of LLMs directly interacting with code; an area with substantial potential but also alot of risk (Yang et al., 2024; Yao et al., 2024a). These models are being deployed in other areas of security already, including writing test cases (Zhang et al., 2023f) and modelling assertions (Kande et al., 2024). Despite their potential utility in defending other systems, the LLMs themselves are highly vulnerable; the vast majority of currently available language models are affected by a new family of attacks, and despite these being specific to LLMs, the defenses are limited (Liu et al., 2023b). The closely related field of adversarial machine learning provides many examples of slight manipulations in a model’s input causing substantial changes in its output (Wiyatno et al., 2019). This holds true for generative models in a similar way, where semantically identical but subtly reworded prompts generate incorrect outputs for robotic tasks (Wu et al., 2024a).
@@
Due to fears of LLMs being used for malicious intent, they generally have “guardrails” integrated (Dong et al., 2024). The exact details of these are not made public, but it is likely that the system message has certain prompts which forbid them to answer questions which are deemed unacceptable. The attack known as “jailbreaking” circumvents these guardrails through intelligent prompting (Greshake et al., 2023). It is difficult to define what constitutes unacceptable inputs to an LLM, and these are often defined by crowd-sourced judgement (Thoppilan et al., 2022). This is closely related to injection attacks which convince the LLM to output privileged information or even run code by accessing a website (Wu et al., 2024b; Liu et al., 2023b). Defenses are currently limited and mostly respond to a specific set of test data with little guarantee that further modification of the attack payload will not defeat the defenses (Chen et al., 2024b). Some approaches employ a conversational approach to finding and preventing security issues, either through conversation with a human (Xu et al., 2021), or between agents (Zeng et al., 2024). Generally, LLMs are often closed-source black box models with unknown architecture. Specific attacks can be used to discover and effectively steal some fraction of the LLM at low cost, which could offer attackers the ability to reverse-engineer compromise the system (Carlini et al., 2024). Other attacks include appending a random set of characters to the end of the prompt cause the LLM to answer a forbidden question (Cao et al., 2023). Some methods check whether the LLM should respond before it is called, catching the problem early (Ruan et al., 2023), whereas others generate the response and check that this is safe to do before taking action. The black-box nature of LLM guardrails can make them difficult to counter, although a promising technique entices the model to explain why it rejects a query, and then uses its own language to invoke a forbidden response (Deng et al., 2023). LLMs can also be fooled by asking leading questions that hint at the wrong answer (Mei et al., 2023). There is an inherent trade-off between safety and helpfulness that must be carefully tuned depending on the application (Zhang et al., 2023i).
@@
Regardless of the details, a vulnerable LLM which is connected to robot or multi-robot system poses a considerable threat if bystanders could (deliberately or accidentally) trick it into performing unsafe or undesired actions. For language-based robots to be effectively deployed, there is considerable need for stronger guarantees of security.
5.1.3. Training Cost
@@
Large foundational models have a very high training cost. This becomes particularly problematic if the environment is non-stationary (the reward changes over time) (de Curtò et al., 2023). More generally, if a lot of training effort is applied to a large model, it is specific to the current environment and goals much the same as in biological organisms (Dawkins, 2017). When the situation changes, which is very likely in robotics, the model may need to be adapted. This issue is not specific to LLMs, and LLMs may have the ability to adapt to the changes more intelligently than RL and other approaches, as their basis of training on diverse text gives them some experience of the implications of novel scenarios (Shek et al., 2023). Future robotic systems may need to be more flexible to changing demands and modular architectures may need to be employed to facilitate this (Liang et al., 2022; Devin et al., 2017). Regardless, without new methods to resolve this, LLMs are likely to be expensive and time-consuming to train.
5.2. Issues with LLMs in Robotic Systems

### v3

@@ 3.4.3 Dynamic Adaptation with Language
Some systems use language as a feedback mechanism with which to adjust an existing model. For example, a functional policy for a robot to walk can take in a textual description of the environment and can use this to modify the behaviour accordingly Shek et al., (2023). This provides a promising means for robots to adjust to unseen conditions through the generality of language models. By invoking the LLM to check each step of a plan, iterative adaptation can be achieved within the model’s architecture. This can include checking each step of the plan one-by-one Raman et al., (2022), or calling on external observations from other sources to update, improve, or complete each step of the plan Xu et al., 2023a . Textual feedback can also be used to catch mistakes when a robot’s plan is converted to a Planning Domain Definition Language (PDDL) and the plan is symbolically checked; error messages can be passed back to the model for the plan to be adjusted accordingly Shirai et al., (2023). This same approach can also be used for parsing error messages from robot execution, which in turn are used to edit generated code in the same way a human might respond to error messages while programming Skreta et al., (2023). Alternatively, simple assessments of why a plan did not work are then converted into summaries and revised plans Shinn et al., (2024). Generally speaking, these methods use the environment (whether that be a physical environment or a human) as a means to evaluate their solution and then iteratively improve the output through this process Yao et al., (2023); Wu et al., 2023c . Another approach is to generate a language plan step by step but to execute the plan as the steps are generated. After a step is completed, an environment update is fed back into the model so it receives real-time feedback and can adjust accordingly Jin et al., (2023). Adaptation can also be used to adapt a shared plan iteratively for different goals or to resolve different issues, for example, a “brain LLM” can write a high-level plan which is adjusted by a “body LLM” to make it performable by the robot. This allows a single plan to be written, improved, and ultimately realised by successive language calls Bhat et al., (2024).
4 Applications
@@ 4.1 Autonomous Vehicles
Although human drivers do not generally have the ability to communicate verbally with each other, some amount of implicit communication is used (pointing, flashing lights, etc). It is possible that the integration of LLMs will provide agents an interface with which to perform the complex coordination required on road environments with a wide variety of vehicles each pursuing different end goals. The changeable nature of road systems such as contraflows, breakdowns, and weather is very difficult to train for with existing approaches, so LLMs may provide a useful means to interpret and handle unexpected scenarios Cui et al., 2024b . A few works are actively pursuing LLMs as a control method for vehicles, taking in textual descriptions and sometimes images, and outputting a text-based command for how the steering wheel and pedals should be adjusted Cui et al., 2023a ; Cui et al., 2024a ; Xu et al., 2023c . A key advantage of using LLMs in this way is that they can provide reasoning and justification, either as an observational tool to explain the actions of an existing model Park et al., (2024), or to pair a reason with the predicted action of the human, which in turn could be used to control the vehicle Wen et al., (2023). The poor ability to reason about the environment, along with obvious safety concerns likely render a fully integrated language-based self-driving car unlikely. Some have addressed this with hybrid approaches, combining the LLM with a more reliable rule-based planner Sharan et al., (2023). However, the most practical avenue for LLMs in this domain is language as a tool for explanation (robot-to-human), control at a very high level (human-to-robot), and inter-agent negotiation e.g. at a stop junction (robot-to-robot). Such negotiation is naturally suited to LLMs, whether in a competitive setting Wu et al., 2024c , or a more cooperative system Hunt et al., (2024). The explainability problem is especially well-matched to autonomous vehicles, and is being actively explored Tekkesinoglu et al., (2024); language is an especially promising avenue in this space and naturally extends the way humans might explain their intentions and actions while driving Yuan et al., 2024a .
4.2 Language-based Explainability of Robot Behaviour
@@
Explainability for language models is a popular area of research with a wide variety of sub-areas; the literature is still young and LLMs are still fundamentally a black-box technology that is difficult to understand Zhao et al., (2024). Despite this, language is a natural domain for explanation, especially for non-technical users. LLMs are themselves used to explain the behaviour of other models González-Santamarta et al., (2023), offering the additional advantage of open dialogue where a human can ask specific targetted questions of the model Lakkaraju et al., (2022). Robotics has an especially high demand for explainability, as robots tend to be mistrusted in general Lewis et al., (2018), and can often cause damage or loss of life if they malfunction Setchi et al., (2020). The union of language and robotics creates an area with an absolute need for explainability and transparency in the absence of trustworthy execution. Robots that use language would be able to explain their failures Das et al., (2021), or could continuously update operators and surrounding humans of their reasoning, intentions, and beliefs. This type of explainability may prove crucial to providing safe and trustworthy AI systems in the future.
4.3 Ad-hoc coordination
@@
As robots and other autonomous systems become more prevalent it will be necessary for teams of such devices to dynamically form and cooperate Stone et al., (2010). For example, several home robots may need to organise a group strategy to clean a home, a team of agricultural robots may need to combine efforts to expedite a task before the weather turns, or a set of helper robots might even need to collaborate to save a human in danger. Each of these requires a coordination and control structure that is agnostic to the models and morphology of the platforms involved Flavin and Sen, (2024). Language is already used by humans in this way and is an important aspect of effective ad-hoc teams White et al., (2018); Mirsky et al., (2020). Collaborative emergent behaviour can be achieved through the use of language Li et al., 2023b . LLM-based structures may also provide a means for differences in robot ability to be considered and optimised for during allocation and planning Han et al., 2024b . In these types of systems, a variety of possible topologies may be required featuring peer-to-peer or hierarchical structures, each requiring an organisational system which could benefit from language as a communication medium in the same way humans operate Guo et al., (2024). In order to establish a shared understanding, robots may need to agree upon names and labels for objects in the environment based on sensor observations, after which point language can be used to represent the complex details of this specific entity when coordinating towards a goal Ekila et al., (2024). Finally, this type of coordination naturally extends to human-robot teams Liu et al., 2023a ; Gong et al., (2023).
4.4 Cobots
@@
The popular emerging model for robots in the workplace is the Collaborative Robot (“cobot”) model. This means that humans are not replaced with machines, but instead, humans and robots work alongside each other. Many governments are pursuing this model to improve their local industries, especially for small-to-medium-sized businesses without the need or budget for large and specialised equipment Lefranc et al., (2022). A key advantage of this is that systems and processes could feasibly be structured to utilise the best of each party’s skills and abilities. Much research and many robot platforms have been applied to this problem Taesi et al., (2023). Use cases of real deployment have been performed with reasonable success El Makrini et al., (2018). Many different approaches have been proposed ranging from working simultaneously on the same widget to forming a human-robot production line with sequential work El Zaatari et al., (2019). Regardless, this means of interacting with robots will likely soon be commonplace in not just the factory setting, but in the service and medical domains too Djuric et al., (2016). Substantial trust issues exist due to the dangerous nature of industrial robots Kopp et al., (2022) which may be abated somewhat through the careful use of language. For humans to effectively work alongside robots there must be a shared understanding and means of communication, at least in one direction as a way to task the robot (human-to-robot) Bergman et al., (2019), and language is likely the most powerful and natural approach Ionescu and Schlund, (2021). The inclusion of language in these robots’ policies in any of the ways described in this work could substantially resolve trust issues and provide an effective means for humans and robots to cooperate in this setting.
5 Challenges
@@
There are still many challenges associated with Language Models, both generally and in robotics. In this section, we discuss some of the primary areas of concern.
5.1 Core Issues with LLMs
@@ 5.1.1 Hallucination
The term “hallucination” refers to foundational models inventing or imagining new information that typically has little to do with the inputs or internal knowledge. This problem is well documented and has persisted despite strong concerns that this may render the models almost unusable Rawte et al., (2023). It has been suggested that due to the fundamental architecture and approach of these models, hallucination is impossible to fully eliminate Xu et al., (2024). The attention has turned to mitigating this problem, primarily through detection and resolution, although approaches are varied Tonmoy et al., (2024). The exact reasons behind hallucination are still not well understood, although some have attempted to analyse and find traces of the problem by investigating hidden states in the model Duan et al., (2024). Some have proposed calling in additional knowledge to resolve the issue Liang et al., 2024b , or considering the likelihood of the generated text (for example in vision-language models, a bird near a tree is a sensible output so is more trustworthy) Zhou et al., (2023). Another approach which is relevant to the robot-to-robot category is to collaborate with other agents; some agents can then abstain from a group decision if required Feng et al., (2024). The most popular area of development is to detect hallucination so the outputs of the model can at least be ignored Chen et al., 2023d ; Leiser et al., (2024); Zhu et al., 2024b . This is essential as, like the closely related adversarial examples in image recognition Wiyatno et al., (2019), hallucination events generally give highly erroneous results without a correspondingly low confidence; the model does not know it is ignorant. This has been partially addressed by pairing the outputs of the model with an associated uncertainty value Kirchhof et al., (2024); Han et al., 2024a .
@@
Hallucination is an issue in all LLM applications, although in robotics it may have more impactful negative consequences. An imagined danger which does not exist or an imagined empty space where there is in fact a human could result in a robot causing considerable harm or loss of life. Evidently for an embodied robot to be deployed in the real world, issues surrounding hallucination must be addressed. At the absolute least, such events should reliably be detected so the robot does not misclassify a human, or attempt to act upon an environment it does not truly understand.
5.1.2 Security
@@
Any control system for a robot can be subjected to cyber-attacks. Hostile actors may seek to gain control of or manipulate the robot in order to disrupt industrial processes or ransom a company by interrupting its operations. Connecting an LLM to one of these systems makes it a target. This is especially true with the increasing ubiquity of LLMs directly interacting with code; an area with substantial potential but also a lot of risk Yang et al., (2024); Yao et al., 2024b . These models are being deployed in other areas of security already, including writing test cases Zhang et al., 2023h and modelling assertions Kande et al., (2024). Despite their potential utility in defending other systems, the LLMs themselves are highly vulnerable; the vast majority of currently available language models are affected by a new family of attacks, and despite these being specific to LLMs, the defenses are limited Liu et al., 2023b . The closely related field of adversarial machine learning provides many examples of slight manipulations in a model’s input causing substantial changes in its output Wiyatno et al., (2019). This holds true for generative models in a similar way, where semantically identical but subtly reworded prompts generate incorrect outputs for robotic tasks Wu et al., 2024b .
@@
Due to fears of LLMs being used for malicious intent, they generally have “guardrails” integrated Dong et al., (2024). The exact details of these are not made public, but it is likely that the system message has certain prompts which forbid them to answer questions which are deemed unacceptable. The attack known as “jailbreaking” circumvents these guardrails through intelligent prompting Greshake et al., (2023). It is difficult to define what constitutes unacceptable inputs to an LLM, and these are often defined by crowd-sourced judgement Thoppilan et al., (2022). This is closely related to injection attacks which convince the LLM to output privileged information or even run code by accessing a website Wu et al., 2024a ; Liu et al., 2023b . Defenses are currently limited and mostly respond to a specific set of test data with little guarantee that further modification of the attack payload will not defeat the defenses Chen et al., 2024a . Some approaches employ a conversational approach to finding and preventing security issues, either through conversation with a human Xu et al., (2021), or between agents Zeng et al., (2024). LLMs are often closed-source black box models with unknown architecture. Specific attacks can be used to discover and effectively steal some fraction of the LLM at low cost, which could offer attackers the ability to reverse-engineer and compromise the system Carlini et al., (2024). Other attacks include appending a random set of characters to the end of the prompt, which causes the LLM to answer a forbidden question Cao et al., (2023). Some methods check whether the LLM should respond before it is called, catching the problem early Ruan et al., (2023), whereas others generate the response and check that this is safe to do before taking action. The black-box nature of LLM guardrails can make them difficult to counter, although a promising technique entices the model to explain why it rejects a query, and then uses its own language to invoke a forbidden response Deng et al., (2023). LLMs can also be fooled by asking leading questions that hint at the wrong answer Mei et al., (2023). There is an inherent trade-off between safety and helpfulness that must be carefully tuned depending on the application Zhang et al., 2023i .
@@
Regardless of the details, a vulnerable LLM which is connected to a robot or multi-robot system poses a considerable threat if bystanders could (deliberately or accidentally) trick it into performing unsafe or undesired actions. For language-based robots to be effectively deployed, there is a considerable need for stronger guarantees of security.
5.1.3 Training Cost
@@
Large foundational models have a very high training cost. This becomes particularly problematic if the environment is non-stationary (the reward or expected state of the environment changes over time) de Curtò et al., (2023). More generally, if a lot of training effort is applied to a large model, it is specific to the current environment and goals much the same as in biological organisms Dawkins, (2017). When the situation changes, which is very likely in robotics, the model may need to be adapted. This issue is not specific to LLMs, and LLMs may have the ability to adapt to the changes more intelligently than RL and other approaches, as their basis of training on diverse text gives them some experience of the implications of novel scenarios Shek et al., (2023). Future robotic systems may need to be more flexible to changing demands, and so modular architectures may need to be employed to facilitate this Liang et al., (2022); Devin et al., (2017). Regardless, without new methods to resolve this, LLMs are likely to be expensive and time-consuming to train.
5.2 Issues with LLMs in Robotic Systems